# Incentivizing Reasoning for Advanced Instruction-Following of Large Language Models

### v1

@@ A.4 DeepScaleR-Preview-Dataset
"problem": "Let $a_{n}=6^{n}+8^{n}$. Determine the remainder upon dividing $a_{83}$ by $49$.",
"solution": "$6^{83}+8^{83}=(6+8)(6^{82}-6^{81}8+\ldots-8^{81}6+8^{82})$
@@
$6^{83}+8^{83}\equiv 14\cdot 6\equiv\boxed{035}\pmod{49}$",
"answer": "35",
@@ A.5.2 Training
For each experiment on 1.5B, 7B, and 8B models, we use the same Qwen2.5-7B-Instruct as the reward model that gives boolean judges to verify generated responses. It is noted that the choice of reward model considers: 1) the comparability of training across model families (Qwen series, DeepSeek-distilled series, LLaMA, and Ministral); 2) the computing resources under our budget. We believe that the larger, stronger reward model (e.g., Qwen2.5-72B-Instruct, DeepSeek V3) would provide more accurate judgement. To evaluate the competence of the Qwen2.5-7B-Instruct as a LLM judge, we perform both automatic and manual analysis on its scorings. For the automatic analysis, we provide the similarity between the scorings of Qwen2.5-7B-Instruct, Qwen2.5-72B-Instruct, QwQ-32B and those of DeepSeek R1 on 1K randomly sampled generations (see Table 8). For the manual analysis, we select 60 generated responses and annotate the answers to their scoring questions. The accuracy of DeepSeek R1, QwQ-32B, Qwen2.5-7B-Instruct, and Qwen2.5-72B-Instruct is reported in Table 9. It can be observed from Table 8 that compared with the DeepSeek R1, the Qwen2.5-7B-Instruct model indeed achieves a high recall that can pinpoints correct responses. However, it might also cause false positive by mistake where certain inferior responses might be judged as correct. From Table 9, we can see that the average accuracy of Qwen2.5-7B-Instruct is around 68.8%, which is slightly lower than Qwen2.5-72B-Instruct. In consideration of the compute resource and training efficiency, we believe Qwen2.5-7B-Instruct is indeed an acceptable "proxy" reward model. Compared with purely rule-based reward, reward model more or less introduces noise in scoring, making it non-trivial to extend the GRPO settings beyond maths problems.
@@
System-level Trigger-Instruction for CoT reasoning You are a helpful assistant. The assistant first thinks about the reasoning process in the mind and then provides the user with the answer. The reasoning process and answer are enclosed within <think> </think> and <answer> </answer> tags, respectively, i.e., <think> reasoning process here </think><answer> answer here </answer>. Now the user asks you to complete a task. After thinking, when you finally reach a conclusion, make sure all your final responses are enclosed between one <answer> tag and one </answer> tag.
@@ A.8 Limitations and Future Work
"https://en.wikipedia.org/wiki/Raymond_III_Count_of_Tripoli". Do not use any commas and highlight at least 3 sections that has titles in markdown format for example *highlighted section part 1* *highlighted section part 2* *highlighted section part 3*.
@@
| 问题 | 关键词 | 简介 | 步骤 |
| 智能清洁机器人 | 智能清洁 | 通过AI算法优化清洁路径,减少人工干预 | 1. 选择适合的智能清洁机器人品牌(如iRobot或Neato)<br>2. 安装机器人并设置清洁区域<br>3. 定期更新地图和清洁计划 |
| 可穿戴健康监测设备 | 健康监测 | 实时监测心率、血压等生理指标,预防疾病 | 1. 选择可穿戴设备(如Apple Watch或Fitbit)<br>2. 连接设备到手机应用<br>3. 定期查看健康数据并调整生活习惯 |
| 番茄工作法时间管理 | 时间管理 | 通过25分钟专注工作+5分钟休息的循环提高效率 | 1. 下载番茄钟应用(如Forest或Pomodoro Timer)<br>2. 设置25分钟工作时间<br>3. 每完成一个番茄钟后休息5分钟 |
| 零废弃生活 | 环保 | 减少一次性塑料使用,实现资源循环利用 | 1. 准备可重复使用的购物袋、餐具和水杯<br>2. 分类回收垃圾,参与社区回收活动<br>3. 购买本地和有机食品,减少运输碳足迹 |
| 智能家居系统 | 智能控制 | 通过语音或手机控制家中的灯光、温度等 | 1. 选择智能音箱(如Amazon Echo或Google Home)<br>2. 连接智能插座、温控器等设备<br>3. 设置场景模式,如“回家模式”或“离家模式” |
@@
Response to an example from ComplexBench Wordsworth’s "I Wondered Lonely as a Cloud" has another name, Growth of a Poet’s Mind.
判断上述陈述的对错,如果正确,则先回答关键词"正确",如果错误,则先回答关键词"错误"。然后陈述原因。
@@
"Avocado vs. Candy Bar: Can You Eat One Daily Without Gaining Weight?"

### v2

@@ A.4 DeepScaleR-Preview-Dataset
”problem”: ”Let $a_{n}=6^{n}+8^{n}$. Determine the remainder upon dividing $a_{83}$ by $49$.”,
”solution”: ”$6^{83}+8^{83}=(6+8)(6^{82}-6^{81}8+\ldots-8^{81}6+8^{82})$
@@
$6^{83}+8^{83}\equiv 14\cdot 6\equiv\boxed{035}\pmod{49}$”,
”answer”: ”35”,
@@ A.5.2 Training
For each experiment on 1.5B, 7B, and 8B models, we use the same Qwen2.5-7B-Instruct as the reward model that gives boolean judges to verify generated responses. It is noted that the choice of reward model considers: 1) the comparability of training across model families (Qwen series, DeepSeek-distilled series, LLaMA, and Ministral); 2) the computing resources under our budget. We believe that the larger, stronger reward model (e.g., Qwen2.5-72B-Instruct, DeepSeek V3) would provide more accurate judgement. To evaluate the competence of the Qwen2.5-7B-Instruct as a LLM judge, we perform both automatic and manual analysis on its scorings. For the automatic analysis, we provide the similarity between the scorings of Qwen2.5-7B-Instruct, Qwen2.5-72B-Instruct, QwQ-32B and those of DeepSeek R1 on 1K randomly sampled generations (see Table 8). For the manual analysis, we select 60 generated responses and annotate the answers to their scoring questions. The accuracy of DeepSeek R1, QwQ-32B, Qwen2.5-7B-Instruct, and Qwen2.5-72B-Instruct is reported in Table 9. It can be observed from Table 8 that compared with the DeepSeek R1, the Qwen2.5-7B-Instruct model indeed achieves a high recall that can pinpoints correct responses. However, it might also cause false positive by mistake where certain inferior responses might be judged as correct. From Table 9, we can see that the average accuracy of Qwen2.5-7B-Instruct is around 68.8%, which is slightly lower than Qwen2.5-72B-Instruct. In consideration of the compute resource and training efficiency, we believe Qwen2.5-7B-Instruct is indeed an acceptable ”proxy” reward model. Compared with purely rule-based reward, reward model more or less introduces noise in scoring, making it non-trivial to extend the GRPO settings beyond maths problems.
@@
System-level Trigger-Instruction for CoT reasoning You are a helpful assistant. The assistant first thinks about the reasoning process in the mind and then provides the user with the answer. The reasoning process and answer are enclosed within ¡think¿ ¡/think¿ and ¡answer¿ ¡/answer¿ tags, respectively, i.e., ¡think¿ reasoning process here ¡/think¿¡answer¿ answer here ¡/answer¿. Now the user asks you to complete a task. After thinking, when you finally reach a conclusion, make sure all your final responses are enclosed between one ¡answer¿ tag and one ¡/answer¿ tag.
@@ A.8 Limitations and Future Work
”https://en.wikipedia.org/wiki/Raymond_III_Count_of_Tripoli”. Do not use any commas and highlight at least 3 sections that has titles in markdown format for example *highlighted section part 1* *highlighted section part 2* *highlighted section part 3*.
@@
— 问题 — 关键词 — 简介 — 步骤 —
— 智能清洁机器人 — 智能清洁 — 通过AI算法优化清洁路径,减少人工干预 — 1. 选择适合的智能清洁机器人品牌(如iRobot或Neato)¡br¿2. 安装机器人并设置清洁区域¡br¿3. 定期更新地图和清洁计划 —
— 可穿戴健康监测设备 — 健康监测 — 实时监测心率、血压等生理指标,预防疾病 — 1. 选择可穿戴设备(如Apple Watch或Fitbit)¡br¿2. 连接设备到手机应用¡br¿3. 定期查看健康数据并调整生活习惯 —
— 番茄工作法时间管理 — 时间管理 — 通过25分钟专注工作+5分钟休息的循环提高效率 — 1. 下载番茄钟应用(如Forest或Pomodoro Timer)¡br¿2. 设置25分钟工作时间¡br¿3. 每完成一个番茄钟后休息5分钟 —
— 零废弃生活 — 环保 — 减少一次性塑料使用,实现资源循环利用 — 1. 准备可重复使用的购物袋、餐具和水杯¡br¿2. 分类回收垃圾,参与社区回收活动¡br¿3. 购买本地和有机食品,减少运输碳足迹 —
— 智能家居系统 — 智能控制 — 通过语音或手机控制家中的灯光、温度等 — 1. 选择智能音箱(如Amazon Echo或Google Home)¡br¿2. 连接智能插座、温控器等设备¡br¿3. 设置场景模式,如“回家模式”或“离家模式” —
@@
Response to an example from ComplexBench Wordsworth’s ”I Wondered Lonely as a Cloud” has another name, Growth of a Poet’s Mind.
判断上述陈述的对错,如果正确,则先回答关键词”正确”,如果错误,则先回答关键词”错误”。然后陈述原因。
@@
”Avocado vs. Candy Bar: Can You Eat One Daily Without Gaining Weight?”